### v1

@@ IV Bethe Ansatz Equations
We now turn to present the exact solution of the Hamiltonian Eq.(1). Since the scattering matrices of the model satisfy Yang-Baxter equations as shown in Appendix C, there is no particle production or annihilation during scattering processes in this theory. Thus, it is possible to study the model by means of the Bethe Ansatz in the presence of volume cut-off $L$, the length of the line segment with a fixed $N$ number of particles. We derive the Bethe Ansatz equations using the fusion hierarchy [71, 68, 65] and Baxter’s T-Q relation [72, 65] in Appendix C in the presence of the volume cutoff $L$ and the UV cutoff $K=2\pi D$ where $D=\frac{N}{L}$ is the particle density.
@@
$\prod_{\upsilon=\pm}\left(\frac{\lambda_{\alpha}+\upsilon d+\frac{i}{2}}{\lambda_{\alpha}+\upsilon d-\frac{i}{2}}\right)\left(\frac{\lambda_{\alpha}+\upsilon b+i}{\lambda_{\alpha}+\upsilon b-i}\right)^{N}=\prod_{\beta\neq\alpha}\frac{\lambda_{\alpha}-\lambda_{\beta}+i}{\lambda_{\alpha}-\lambda_{\beta}-i}\frac{\lambda_{\alpha}+\lambda_{\beta}+i}{\lambda_{\alpha}+\lambda_{\beta}-i},$ (8)
@@
The problem of diagonalizing the Hamiltonian Eq.(1) requires consistently solving the coupled algebraic equations Eq.(7) and Eq.(8). The first term in Eq.(8), which depends on the RG invariant parameter $d$, encodes all the information about the boundary physics. As we shall see, for various values of the parameter $d$, this term gives different contributions to a physical quantity, and hence, this gives us a way to describe the physics of the model in terms of the RG invariant parameter $d$, which could take real or imaginary value. We shall describe the boundary physics in four distinct phases: the Kondo phase, the zero mode phase, the YSR phase, and the unscreened phase with decreasing amount of details.
@@ V The Overscreened Kondo phase
When the RG invariant $d$ is real or imaginary $d=i\delta$ with $10<\delta<$ 1, the impurity is overscreened by the Kondo cloud.
The ground state of the model is described by a sea of two strings solutions $\lambda_{\alpha}=\chi_{\alpha}\pm\frac{i}{2}$ [33] and a hole which corresponds to, we will see, to spin-1/2 propagating spinon in the ground state. It is required for consistency reasons and counting of the states but does not participate in the impurity dynamics, as we shall briefly discuss later.
The Bethe equations for the centers of the two strings $\chi_{\alpha}$ reads
$\prod_{\nu=\pm}\left(\frac{\chi_{\alpha}+\nu d+i}{\chi_{\alpha}+\nu d-i}\right)\left(\frac{\chi_{\alpha}+\nu b+\frac{3i}{2}}{\chi_{\alpha}+\nu b-\frac{3i}{2}}\frac{\chi_{\alpha}+\nu b+\frac{i}{2}}{\chi_{\alpha}+\nu b-\frac{i}{2}}\right)^{N}\frac{\chi_{\alpha}+i}{\chi_{\alpha}-i}=\prod_{\nu=\pm}\prod_{\beta}^{M}\frac{\left(\chi_{\alpha}+\nu\chi_{\beta}+i\right)^{2}}{\left(\chi_{\alpha}+\nu\chi_{\beta}-i\right)^{2}}\frac{\chi_{\alpha}+\nu\chi_{\beta}+2i}{\chi_{\alpha}+\nu\chi_{\beta}-2i}$ (10)
@@
Since we are interested in thermodynamics limit $N\to\infty$, the higher order terms are negligible. Here,
@@ V.1 Excitations
In the Kondo regime (i.e., when $d\in R$ or $d=i\delta$ such that $0<\delta<\frac{1}{2}$ shown in the phase diagram Fig.2), there are no boundary excitations; all the excitations are bulk excitations. A thorough discussion of the bulk excitation can be found at [33]; here, we shall briefly discuss them. The simplest excitations involve spinless charge excitations.
@@
Chargeless spin exciations- are constructed by changing the spin quantum numbers $I_{\gamma}$. For example, elementary spin excitation can be constructed by adding a pair of holes in the sea of two strings in position $\theta_{1}$ and $\theta_{2}$ and a one-string solution at position $\lambda$. The change in the root density is given by
@@
and the energy of this excitation is give
@@
We shall skip the details presented in [33] and describe other fundamental excitations. By creating two holes in the sea of two-strings and adding one one-string and one three-string solution, we obtain singlet excitation with spin $S^{z}_{S}=0$ and energy that is same as that of triplets in the thermodynamic limit i.e.
@@
All other excitations can be constructed by adding more bulk string solutions, even the number of holes, quartets solutions, etc [73].
@@ V.2 Density of states
From the root density Eq.(19), we compute the density of state using relation $\rho_{\mathrm{DOS}}(E)=2\left|\rho_{\mathrm{gs}}(\chi)/E^{\prime}(\chi)\right|$ where $\rho_{gs}(\chi)$ is the inverse Fourier transform of the ground state root density given in Eq.(19). Notice that the root density naturally separates into $\rho_{\mathrm{gs}}(\chi)=N\rho_{\mathrm{bulk}}(\chi)+\rho_{\mathrm{imp}}(\chi,d)+\rho_{\mathrm{boundary}}(\chi)$.
@@
The impurity Kondo phase is characterized by a dynamically generated, RG invariant energy scale $T_{K}$ characterizing its response to external probes such as temperatures of magnetic fields. It is obtained via
@@
Figure 4: The plot of $f(d)=T_{K}/m$ as a function of the RG invariant parameter $d$ in the Kondo and the zero mode phase. Notice that $T_{K}\geq m$ for all values of $d$ and the minimum $T_{K}=m$ occurs at $d=i\delta=i$, which is the boundary between the zero mode phase and the YSR phase.
@@ VI The zero mode Phase
When the impurity parameter $d$ takes the purely imaginary value $d=i\delta$ where $\frac{1}{2}<\delta<1$, the impurity is in the zero mode phase as shown in the phase diagram Fig.2. In this regime, the ground state is composed of the 2-string solutions and a hole with the root density given by
@@
which is due to there being a spin-$\frac{1}{2}$ propagating spinon in the ground state. The ground state is just an analytic continuation of the Kondo phase and is therefore characterized by the dynamically generated scale $T_{K}=mf(\delta)$. The bulk excitations are also the same as in the Kondo phase, which is constructed by adding even numbers of holes and bulk strings as described in Sec.V.1. The impurity density of states in the ground state is the analytic continuation of the density of states in the Kondo phase given by Eq.(31) to imaginary $d=i\delta$.
However, there exists a unique boundary excitation in this phase, where a boundary-bound mode with zero energy exists. This mode is described by the purely imaginary solution of the Bethe Ansatz equations Eq.(8). More precisely, when the RG invariant parameter $d=i\delta$ is purely imaginary and $\delta>\frac{1}{2}$, the Bethe equation Eq.(8) has a purely imaginary solution
@@
The purely imaginary Bethe roots, associated with bound states, appear as poles in the physical or dressed boundary scattering matrix [74, 75, 76, 77]. These boundary string solutions describing the boundary excitation exist in various integrable systems with boundaries[78, 27, 79, 1, 2, 79, 80, 3, 81, 82].
@@
such that the total energy contribution due to this solution obtained from Eq.(12) is exactly equal but with the opposite sign to that of its bare energy $D\tan^{-1}\left(\frac{b}{\frac{3}{2}-\delta}\right)+D\tan^{-1}\left(\frac{b}{\delta+\frac{1}{2}}\right)$. Hence, the total energy of this vanishes in the thermodynamic limit.
The state that is degenerate to the ground state in the thermodynamic limit can thus be constructed by adding a boundary string solution and a one-string at position $\lambda_{c}$, which has the root density of the following form
@@
The spin of this state is $S^{z}=\pm\frac{1}{2}$ (carried by the spinon), which is the same as that of the ground state because the one-string at position $\lambda_{c}$ and the boundary string $\lambda_{\alpha}$ form a singlet. Since both the boundary string and the one-string have vanishing energy in the thermodynamic limit, this two-fold state degenerates with the ground state.
The position of the one string at $\lambda_{c}$ is fixed by the Bethe Ansatz equation written for this root which takes the form
@@
When the free parameter $\theta$ is set to 00 to minimize the energy, the position of $\lambda_{c}$ takes a simple value $\lambda_{c}=\frac{\sinh^{-1}\left(\sqrt{\cosh(\pi\delta)}\right)}{\pi}$.
Notice that even though the ground state and the zero energy excited state are degenerate in the thermodynamic limit, these are physically quite distinct states. As mentioned earlier, in the ground state, the density of the state is the analytic continuation of that of the Kondo ground state. In the zero energy excited state, the density of states in the zero energy excited state contribution from the impurity, one string, and the boundary string is given by
@@
Here, the first term is the analytic continuation of the impurity contribution in the Kondo phase, the second term is the contribution from the boundary string solution, the third term is a contribution from the one string, and the delta function is from the zero energy states. Notice that the integrated density of state contribution from each the string solution and one string is $-\frac{1}{2}$, which is exactly canceled by the delta function term, and the other first term has integrated density of state contribution of $\frac{1}{4}$, which shows that the impurity is screened in this state.
@@ VII The YSR Phase
As mentioned earlier, when $\delta<1$, the impurity is screened in the ground state by a multiparticle cloud of massive spinons, but at $\delta=1$, the model undergoes boundary quantum phase transition, and for $\delta>1$, the impurity is no longer screened in the ground state. In this section, we shall discuss the regime $21<\delta<$ 2 dubbed as the YSR phase in the phase diagram Fig.2. In this phase, the boundary string solution described in Eq.(37) acquires mass and gives rise to a mid-gap state above the ground state akin to the YSR state in the BCS superconductor, and in this mid-gap state, the impurity is screened  [^3] Notice that this is in contrast with the model with spin-$\frac{1}{2}$ bulk studied in [2] where the YSR state is the ground state for some parametric regime – here the YSR regime is a mid-gap excited state throughout the parametric regime $21<\delta<$ 2. As we shall show below, this state requires an additional hole on top of this boundary string solution such that the energy of this state is below the massgap. In the next section, we’ll discuss the region $\delta>2$ where the impurity cannot be screened at any energy scale, though the boundary string solution exists and is gapless. But because the boundary sting becomes a wide string [73], this solution can not be added to the ground state without adding additional massive holes to form an excitation. Such an excitation would, therefore, be above the massgap.
@@
The total number of Bethe roots is $M=2+2\tilde{\rho}_{d}(0)=N,$ and hence, the total magnetization is $S^{z}=N+1/2-M=\frac{1}{2}$, which is the spin of the propagating spinon. Due to $SU(2)$ symmetry, there is another state with $S^{z}=-\frac{1}{2}$. The impurity is screened by this exponentially boundary-localized bound mode in this state.
@@
which is zero at $\delta=1$, attains a maximum value of $2m$ at $\delta=2$, and is monotonically increasing for $21<\delta<$ 2. We shall refer to this excited state as Yu-Shiba-Rushinov (YSR) as the state is akin to the mid-gap Shiba states formed in the BCS superconductor in the presence of classical impurities [25, 24, 26].

### v2

@@ IV Bethe Ansatz Equations
We now turn to present the exact solution of the Hamiltonian Eq.(1). Since the scattering matrices of the model satisfy Yang-Baxter equations as shown in Appendix C, there is no particle production or annihilation during scattering processes in this theory. Thus, it is possible to study the model by means of the Bethe Ansatz in the presence of a volume cut-off $L$, the length of the line segment with a fixed $N$ number of particles. We derive the Bethe Ansatz equations using the fusion hierarchy [71, 69, 66] and Baxter’s T-Q relation [72, 66] in Appendix C in the presence of the volume cutoff $L$ and the UV cutoff $K=2\pi D$ where $D=\frac{N}{L}$ is the particle density.
@@
$\prod_{\upsilon=\pm}\left(\frac{\lambda_{\alpha}+\upsilon d+\frac{i}{2}}{\lambda_{\alpha}+\upsilon d-\frac{i}{2}}\right)\left(\frac{\lambda_{\alpha}+\upsilon b+i}{\lambda_{\alpha}+\upsilon b-i}\right)^{N}$
$=\prod_{\beta\neq\alpha}\frac{\lambda_{\alpha}-\lambda_{\beta}+i}{\lambda_{\alpha}-\lambda_{\beta}-i}\frac{\lambda_{\alpha}+\lambda_{\beta}+i}{\lambda_{\alpha}+\lambda_{\beta}-i},$ (8)
@@
Having solved for the momenta $k_{j}$, the energy of an eigenstate is then given by
$E=\sum_{j}k_{j}.$ (10)
The problem of diagonalizing the Hamiltonian Eq.(1) thus requires consistently solving the coupled algebraic equations Eq.(7) and Eq.(8). The first equation, Eq.(7) is a quantization condition for the pseudomomenta $k_{j}$ (with $L$ being the system size). The second equation (8) constrains the spin rapidities $\lambda_{\alpha}$. Every configuration of $\lambda_{\alpha}$ that satisfies this equation corresponds to an eigenstate of the system, with its energy determined via the first equation. The spin of the state is obtained by knowing the total number of spin flips $M$ using the relation $S^{z}=\left(N+\frac{1}{2}-2M\right)=\frac{1}{2}$.
In essence, the first equation governs the charge degrees of freedom by fixing the allowed values of $k_{j}$, while the second equation encodes the interactions among the spin degrees of freedom. Together, they capture the full spectrum of the model, with the interplay between charge and spin excitations determining both the eigenstates and their corresponding eigenenergies.
The first term in Eq.(8), which depends on the RG invariant parameter $d$, encodes all the information about the boundary physics. As we shall see, for various values of the parameter $d$, this term gives different contributions to a physical quantity, and hence, this gives us a way to describe the physics of the model in terms of the RG invariant parameter $d$, which could take real or imaginary values. We shall describe the boundary physics in four distinct phases: the Kondo phase, the zero mode phase, the YSR phase, and the unscreened phase with decreasing amount of details.
@@ V The Overscreened Kondo phase
When the RG invariant $d$ is real or imaginary $d=i\delta$ with $10<\delta<$ 1, the impurity is overscreened by the Kondo cloud. The ground state of the model consists of a sea of two-string solutions $\lambda_{\alpha}=\chi_{\alpha}\pm\frac{i}{2}$ [33] and a hole which corresponds to a spin-1/2 propagating spinon in the ground state. The spinon is required for consistency reasons and counting of the states but does not participate in the impurity dynamics, as we shall briefly discuss later.
The Bethe equations for the centers of the two strings $\chi_{\alpha}$ take the form
$\prod_{\nu=\pm}\left(\frac{\chi_{\alpha}+\nu d+i}{\chi_{\alpha}+\nu d-i}\right)\left(\frac{\chi_{\alpha}+\nu b+\frac{3i}{2}}{\chi_{\alpha}+\nu b-\frac{3i}{2}}\frac{\chi_{\alpha}+\nu b+\frac{i}{2}}{\chi_{\alpha}+\nu b-\frac{i}{2}}\right)^{N}\frac{\chi_{\alpha}+i}{\chi_{\alpha}-i}=\prod_{\nu=\pm}\prod_{\beta}^{M}\frac{\left(\chi_{\alpha}+\nu\chi_{\beta}+i\right)^{2}}{\left(\chi_{\alpha}+\nu\chi_{\beta}-i\right)^{2}}\frac{\chi_{\alpha}+\nu\chi_{\beta}+2i}{\chi_{\alpha}+\nu\chi_{\beta}-2i}.$ (11)
@@
Since we are interested in the thermodynamic limit $N\to\infty$, the higher order terms are negligible. Here,
@@ V.1 Excitations
In the Kondo regime (i.e., when $d\in R$ or $d=i\delta$ such that $0<\delta<\frac{1}{2}$ shown in the phase diagram Fig.2), there are no boundary excitations; all the excitations are bulk excitations. A thorough discussion of the bulk excitation can be found at [33]; here, we shall briefly discuss it. The simplest excitations involve spinless charge excitations.
@@
Spin exciations - are constructed by changing the spin quantum numbers $I_{\gamma}$. For example, elementary spin excitation can be constructed by adding a pair of holes in the sea of two strings in position $\theta_{1}$ and $\theta_{2}$ and a one-string solution at position $\lambda$. The change in the root density is given by
@@
and the energy of this excitation is given
@@
We shall skip the details presented in [33] and describe other fundamental excitations. By creating two holes in the sea of two-strings and adding one one-string and one three-string solution, we obtain a singlet excitation with spin $S^{z}_{S}=0$ and energy that is the same as that of triplets in the thermodynamic limit i.e.
@@
All other excitations can be constructed by adding more bulk string solutions, even the number of holes, quartet solutions, etc. [73].
@@ V.2 Density of states
From the root density in in Eq.(20), we compute the density of states using the relation $\rho_{\mathrm{DOS}}(E)=2\left|\rho_{\mathrm{gs}}(\chi)/E^{\prime}(\chi)\right|$ where $\rho_{gs}(\chi)$ is the inverse Fourier transform root density in the ground state given in Eq.(20). Notice that the root density naturally separates into $\rho_{\mathrm{gs}}(\chi)=N\rho_{\mathrm{bulk}}(\chi)+\rho_{\mathrm{imp}}(\chi,d)+\rho_{\mathrm{boundary}}(\chi)$.
@@
The impurity Kondo phase is characterized by a dynamically generated, RG invariant energy scale $T_{K}$ characterizing its response to external probes such as temperatures or magnetic fields. It is obtained via
@@
Figure 4: The plot of $f(d)=T_{K}/m$ as a function of the RG invariant parameter $d$ in the Kondo and the zero mode phase. Notice that $T_{K}\geq m$ for all values of $d$, and the minimum $T_{K}=m$ occurs at $d=i\delta=i$, which is the boundary between the zero mode phase and the YSR phase.
In summary, the overscreened Kondo phase is characterized by the many-body overscreening of the spin-$\frac{1}{2}$ impurity by the strongly interacting bath of spin-1 electrons. This leads to an emergent non-Fermi liquid fixed point, similar to the two-channel Kondo effect, where the impurity entropy at low temperatures takes the value $\frac{1}{2}\ln 2$ as we shall see below. A dynamically generated Kondo scale $T_{K}$, distinct from the bulk superconducting mass gap $\Delta$, governs the screening process. In the bulk, the attractive interaction between fermions leads to the formation of a superconducting order with a dynamically generated mass gap $\Delta=2m$. This boundary phase is stable for sufficiently large Kondo coupling $J$, where the boundary interaction dominates over the bulk superconducting correlations, allowing the impurity to be overscreened by a multiparticle Kondo cloud.
@@ VI The zero mode Phase
When the impurity parameter $d$ takes the purely imaginary value $d=i\delta$ where $\frac{1}{2}<\delta<1$, the impurity is in the zero mode phase as shown in the phase diagram Fig.2. In this regime, the ground state is composed of 2-strings and a hole, with the root density given by
@@
which is due to there being a spin-$\frac{1}{2}$ propagating spinon in the ground state. The ground state is an analytic continuation of the Kondo phase and is characterized by the dynamically generated scale $T_{K}=mf(\delta)$. The bulk excitations are also the same as in the Kondo phase, and are constructed by adding even numbers of holes in bulk strings as described in Sec.V.1. The impurity density of states in the ground state is the analytic continuation of the density of states in the Kondo phase given by Eq.(32) to imaginary $d=i\delta$.
However, there also exists a unique boundary excitation in this phase - a boundary-bound mode with zero energy. This mode is described by the purely imaginary solution of the Bethe Ansatz equations Eq.(8). More precisely, when the RG invariant parameter $d=i\delta$ is purely imaginary and $\delta>\frac{1}{2}$, the Bethe equation Eq.(8) has a purely imaginary solution
@@
The purely imaginary Bethe roots, associated with bound states, appear as poles in the physical or dressed boundary scattering matrix [74, 75, 76, 77]. This kind of boundary string solutions describing the boundary excitation exist in various integrable systems with boundaries[78, 27, 79, 1, 2, 79, 80, 3, 81, 82].
@@
such that the total energy contribution due to this solution obtained from Eq.(13) is exactly equal but with the opposite sign to that of its bare energy $D\tan^{-1}\left(\frac{b}{\frac{3}{2}-\delta}\right)+D\tan^{-1}\left(\frac{b}{\delta+\frac{1}{2}}\right)$. Hence, the total energy of this excitation vanishes in the thermodynamic limit.
The state is degenerate with the ground state in the thermodynamic limit is constructed by adding a boundary string solution and a one-string at position $\lambda_{c}$, resulting in a root density of the following form
@@
The spin of this state is $S^{z}=\pm\frac{1}{2}$ (carried by the spinon), which is the same as that of the ground state because the one-string at position $\lambda_{c}$ and the boundary string $\lambda_{\alpha}$ form a singlet. Since both the boundary string and the one-string have vanishing energy in the thermodynamic limit, this two-fold state is degenerate with the ground state, as mentioned above.
The position of the one string at $\lambda_{c}$ is fixed by the Bethe Ansatz equation written for this root, which takes the form
@@
Notice that even though the ground state and the zero-energy excited state are degenerate in the thermodynamic limit, these are physically quite distinct states. As mentioned earlier, in the ground state, the density of the state is the analytic continuation of that of the Kondo ground state. In the zero energy excited state, the density of states in the zero energy excited state contribution from the impurity, one string, and the boundary string is given by
@@
Here, the first term is the analytic continuation of the impurity contribution in the Kondo phase, the second term is the contribution from the boundary string solution, the third term is a contribution from the one string, and the delta function is from the zero-energy states. Notice that the integrated density of state contribution from each the string solution and one string is $-\frac{1}{2}$, which is exactly canceled by the delta function term, and the other first term has integrated density of state contribution of $\frac{1}{4}$, which shows that the impurity is screened in this state.
In summary, the zero mode phase is an intermediate phase that exhibits characteristics of both the Kondo phase and the YSR phase. As in the former, the impurity is many-body screened by a multiparticle Kondo cloud, while in the latter, there exists a boundary excitation, albeit with vanishing energy in the thermodynamic limit. As the system parameters are tuned such that bulk superconducting strength competes with the Kondo screening mechanism, this phase emerges, retaining the overscreening mechanism of the Kondo phase but introducing an additional boundary excitation. The phase is characterized by the presence of a purely imaginary Bethe root solution, corresponding to a zero-energy localized mode at the boundary. The presence of this additional zero-energy excitation differentiates this phase from the conventional overscreened Kondo regime, marking a novel impurity behavior that arises due to the interplay of bulk and boundary interactions.
@@ VII The YSR Phase
As mentioned earlier, when $\delta<1$, the impurity is screened in the ground state by a multiparticle cloud of massive spinons, but at $\delta=1$, the model undergoes a boundary quantum phase transition, and for $\delta>1$, the impurity is no longer screened in the ground state. In this section, we shall discuss the regime $21<\delta<$ 2 dubbed as the YSR phase in the phase diagram Fig.2. In this phase, the boundary string solution described in Eq.(38) acquires mass and gives rise to a mid-gap state above the ground state akin to the YSR state in the BCS superconductor, and in this mid-gap state, the impurity is screened  [^2] Notice that this is in contrast with the model with spin-$\frac{1}{2}$ bulk studied in [2] where the YSR state is the ground state for some range of parameters – here the YSR regime is a mid-gap excited state throughout the parametric regime $21<\delta<$ 2. As we shall show below, this state requires an additional hole on top of this boundary string solution such that the energy of this state is below the mass gap. In the next section, we shall discuss the region $\delta>2$ where the impurity cannot be screened at any energy scale, though the boundary string solution exists and is gapless. But because the boundary string becomes a wide string [73], this solution cannot be added to the ground state without adding additional massive holes to form an excitation. Such an excitation would, therefore, be above the mass gap.
@@
The total number of Bethe roots is $M=2+2\tilde{\rho}_{d}(0)=N,$ and hence, the total magnetization is $S^{z}=N+1/2-M=\frac{1}{2}$, which is the spin of the propagating spinon? Due to $SU(2)$ symmetry, there is another state with $S^{z}=-\frac{1}{2}$. The impurity is screened by this exponentially boundary-localized bound mode in this state.
@@
which is zero at $\delta=1$, attains a maximum value of $2m$ at $\delta=2$, and is monotonically increasing for $21<\delta<$ 2. We shall refer to this excited state as Yu-Shiba-Rusinov (YSR), as the state is akin to the mid-gap Shiba states formed in the BCS superconductor in the presence of classical impurities [25, 24, 26].